Circular No. 6153

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995I IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery

of an apparent supernova of mag V ∼ 19.5 on two films taken by Pollas and D. Albanese with the OCA 0.9-m Schmidt telescope on Feb. 28.15 and Mar. 1.15 UT. SN 1995I is located at α = 13^{h}20^{m}59^{s}.85, δ = +3^{o}35′56′′.2 (equinox 2000.0), which is 0′′.5 west and 5′′.8 north of the galaxy’s center, at the northern tip of the host galaxy. Another technical pan film (limiting mag 22) taken by Albanese and A. Maury on Mar. 24.08 shows that the object has faded by ∼ 1 mag or more, to mag 20.5-21.0. No object is present at this location on the original Palomar Sky Survey prints. A nearby star (V ∼ 15.2) is at position end figures 21^{m}01^{s}.33, 37′37′′.4.

GX 1+4 D. Chakrabarty, T. Koh, T. A. Prince, and B. Vaughan, California

Institute of Technology; M. H. Finger and M. Scott, Universities Space Research Association; and R. B. Wilson, Marshall Space Flight Center, NASA, report for the Compton Gamma Ray Observatory BATSE team: “The accreting x-ray pulsar GX 1+4 has undergone another torque reversal, and resumed spinning down on around Mar. 2, thereby ending an extended spin-up episode that began on 1994 Oct. 30 (IAUC 6105). On Mar. 17.0 UT, the barycentric pulse frequency was (8.21119 ± 0.00002) × 10^{−3} s^{−1}, with a mean frequency derivative of (−1.25 ± 0.04) × 10^{−12} s^{−2}. The phaseaveraged pulsed intensity on this date was (9.0 ± 0.9) × 10^{−4} photon cm^{−2} s^{−1} keV^{−1} at 30 keV, with a power-law photon index of 2.6 ± 0.2 for the 20- to 100-keV pulsed flux.”

SUPERNOVA 1995F IN NGC 2726 Additional CCD photometry by M. Blake, T. Fuller, D. J. Lane, J.

Asbell-Clarke, S. Mitchell, and S. Elieff, St. Mary’s University (cf. IAUC 6138; comparison star GSC 3811.0095): Feb. 19.01 UT, R = 14.8 ± 0.1; 28.226, 15.2 ± 0.1; Mar. 11.989, 15.2 ± 0.1; 20.021, 15.2 ± 0.2.

SUPERNOVA 1995D IN NGC 2962 Additional CCD photometry (cf. IAUC 6149): Mar. 11.94 UT, V =

14.23 ± 0.03, B − V = +0.79 ± 0.05 (H. Mikuž, Ljubljana); 12.019, R = 14.2 ± 0.1 (Blake et al.); 20.069, R = 14.4 ± 0.1 (Blake et al.).

1995 March 27 Daniel W. E. Green